# Decoupled Classifier-Free Guidance for Counterfactual Diffusion Models

### v2

@@ Abstract
Counterfactual image generation aims to simulate realistic visual outcomes under specific causal interventions. Diffusion models have recently emerged as a powerful tool for this task, combining DDIM inversion with conditional generation via classifier-free guidance (CFG). However, standard CFG applies a single global weight across all conditioning variables, which can lead to poor identity preservation and spurious attribute changes—a phenomenon known as attribute amplification. To address this, we propose Decoupled Classifier-Free Guidance (DCFG), a flexible and model-agnostic framework that introduces group-wise conditioning control. DCFG builds on an attribute-split embedding strategy that disentangles semantic inputs, enabling selective guidance on user-defined attribute groups. For counterfactual generation, we partition attributes into intervened and invariant sets based on a causal graph and apply distinct guidance to each. Experiments on CelebA-HQ, MIMIC-CXR, and EMBED show that DCFG improves intervention fidelity, mitigates unintended changes, and enhances reversibility, enabling more faithful and interpretable counterfactual image generation.
@@ 1 Introduction
Counterfactual generation is fundamental to causal reasoning [37, 40, 1], allowing us to explore hypothetical scenarios, such as How would this patient’s disease have progressed if treatment A had been administered instead of treatment B? The ability to answer such causal questions is important across various domains, such as healthcare [46], fairness [26, 65] and scientific discovery [33]. There has been a growing interest in generating counterfactual images using deep generative models [23, 67], aiming to simulate how visual data would change under hypothetical interventions. Recent works extend the framework of Structural Causal Models (SCMs) [38, 40, 2] with deep generative models, performing counterfactual generation via a three-step procedure: abduction-action-prediction [36, 9, 62, 61, 8, 47]. These works typically parameterise their image-generating mechanisms with normalising flows [35, 59], GANs [14], VAEs [22], and HVAEs [5, 56]. However, these generative backbones often exhibit trade-offs between fidelity and flexibility [43, 64]. In this work, we explore diffusion models as a more powerful alternative for counterfactual inference.
@@
Diffusion models [50, 54, 17] have emerged as the state-of-the-art approach for image synthesis, achieving unprecedented fidelity and perceptual quality [10, 41]. Recent works have utilised diffusion models for counterfactual generation [34, 24, 46, 45, 3, 4, 29, 58, 44, 12, 53, 39, 25]. A common approach involves DDIM inversion to encode observed images into latent representations, followed by conditional generation under modified attributes. Conditioning is typically enforced via discriminative score functions, either with external classifiers [10] or through classifier-free guidance [16]. This combination of DDIM inversion and conditional decoding has become a dominant paradigm in diffusion-based image editing [7, 31, 57, 15, 11]. Building upon this paradigm, we investigate its use for high-fidelity counterfactual inference via a novel classifier-free guidance strategy.
Classifier-Free Guidance (CFG) [16] has become the standard approach for conditioning diffusion models without requiring external classifiers. In counterfactual generation, CFG plays a crucial role in ensuring that the intended intervention is faithfully reflected in the output. While recent works have proposed refinements to CFG to enhance sample fidelity and controllability [6, 21, 27], standard CFG applies a single global guidance weight across all conditioning attributes, regardless of which ones are intended to change. This can lead to over-editing or unintended modifications of attributes that should remain fixed—a phenomenon known as attribute amplification [63], where invariant factors are spuriously exaggerated. While Xia et al. [63] observed this issue in HVAE-based counterfactual models [9] as a result of training-time entanglement, we show that a similar failure mode can arise at inference time in diffusion models due to the indiscriminate application of global guidance. Such behaviour not only violates the underlying causal assumptions—by modifying attributes outside the intervention—but can also cause the generation trajectory to drift away from the original data manifold, degrading identity preservation [66, 31, 55].
@@
To address the limitations of standard CFG, we propose Decoupled Classifier-Free Guidance (DCFG), a model-agnostic method that requires no changes to the diffusion model or loss function, and only modifies the conditioning embedding. DCFG can be readily applied to any diffusion model employing CFG. A key ingredient in our approach is an attribute-split conditioning embedding strategy, which disentangles semantic attributes in the embedding space and enables selective masking and group-wise modulation at inference time. Unlike standard CFG—which applies a single global guidance weight across all conditioning attributes—DCFG assigns separate weights to attribute groups, allowing for fine-grained, interpretable control over the generative process. While conceptually related to compositional diffusion approaches, our method differs significantly: Shen et al. [49] apply pixel-wise spatial masks to modulate guidance locally, and Liu et al. [28] rely on multiple conditional diffusion models fused via shared score functions. In contrast, DCFG uses a single model and modulates guidance at the semantic attribute level. For counterfactual generation, we instantiate DCFG by partitioning attributes into intervened and invariant sets based on their roles in a structural causal graph, and apply distinct guidance to each group. Crucially, by decoupling guidance and focusing it solely on the intended intervention, DCFG reduces the risk of the generation trajectory drifting away from the original data manifold [66, 31, 55]. While our experiments focus on causal groupings, the DCFG framework is general and supports arbitrary partitions of semantic attributes. The contributions of this paper are:
  • We introduce an attribute-split embedding strategy that disentangles semantic conditioning variables in the embedding space, enabling selective masking and modular guidance for DCFG.
@@
  • We propose Decoupled Classifier-Free Guidance (DCFG), a simple and flexible extension of CFG that assigns separate guidance weights to groups of attributes. DCFG is model-agnostic, training-compatible with existing pipelines, and supports arbitrary attribute groupings at inference time.
  • We instantiate DCFG for counterfactual image generation by grouping attributes into intervened and invariant sets based on a structural causal graph, and show improved alignment with intervention targets while preserving irrelevant features.
@@ 2 Background
2.1 Causality
Structural Causal Models. SCMs [37] consist of a triplet $\langle U,A,F\rangle$, where $U=\{u_{i}\}_{i=1}^{K}$ denotes the set of exogenous (latent) variables, $A=\{a_{i}\}_{i=1}^{K}$ the set of endogenous (observed) variables, and $F=\{f_{i}\}_{i=1}^{K}$ a collection of structural assignments such that each variable $a_{k}$ is determined by a function $f_{k}$ of its parents $\mathbf{pa}_{k}\subseteq A\setminus a_{k}$ and its corresponding noise $u_{k}$: $a_{k}\coloneqq f_{k}(\mathbf{pa}_{k},u_{k})$. SCMs enable causal reasoning and interventions via the do-operator, e.g., setting a variable $a_{k}$ to a fixed value $c$ through $do(a_{k}\coloneqq c)$. In this work, we focus on generating image counterfactuals and implement the underlying image synthesis mechanism using diffusion models.
Counterfactual inference. We define an image as $\mathbf{x}$, which is generated from its causal parents $\mathbf{pa}$ and an exogenous noise variable $\mathbf{u}$ via a structural assignment: $\mathbf{x}:=f(\mathbf{u},\mathbf{pa})$. Counterfactual reasoning proceeds in three conceptual steps [37] - (1) Abduction: infer the latent noise $\mathbf{u}$ from the observed data and its parents: $\mathbf{u}:=f^{-1}(\mathbf{x},\mathbf{pa})$; (2) Action: apply an intervention to alter selected parent variables, yielding $\widetilde{\mathbf{pa}}$; (3) Prediction: propagate the effect of the intervention through the model to compute the counterfactual $\tilde{\mathbf{x}}:=f(\mathbf{u},\widetilde{\mathbf{pa}})$. Recent advancements have sought to implement these steps using deep learning models, with VAE [36], HVAE [9], diffusion models [44, 45, 24, 12]. The general idea is to model the structural assignment $f_{\theta}$ and its inverse $f^{-1}_{\theta}$ using neural networks with trainable parameters $\theta$.
2.2 Diffusion Models
DMs [51, 17] are latent variable models which are trained to generate images by gradually removing Gaussian noise from $x_{T}\sim N(0,I)$ over $T$ timesteps. Given a clean sample $\mathbf{x}_{0}\sim p_{0}(\mathbf{x})$, the forward process at timestep $t$ follows:
$\mathbf{x}_{t}=\sqrt{\alpha_{t}}\,\mathbf{x}_{0}+\sqrt{1-\alpha_{t}}\,\boldsymbol{\epsilon},\quad\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (1)
where $\{\alpha_{t}\}_{t=1}^{T}$ is a variance schedule defined over $t\in[1,T]$.
To learn the reverse process, a parameterized network $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})$ is trained to predict the added noise from noisy inputs. We adopt the conditional diffusion model formulation, where $\mathbf{c}$ denotes the embedding of semantic parent attributes $\mathbf{pa}$ used as conditioning input. The training objective minimizes the noise prediction loss:
@@ 2.2 Diffusion Models
At inference time, data samples are generated by progressively denoising $\mathbf{x}_{t}$. Following Eq (12) in Song et al. [52], the denoising step from $\mathbf{x}_{t}$ to $\mathbf{x}_{t-1}$ is:
$\mathbf{x}_{t-1}=\underbrace{\sqrt{\alpha_{t-1}}\left(\frac{\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\,\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})}{\sqrt{\alpha_{t}}}\right)}_{\text{predicted }\mathbf{x}_{0}\text{}}+\underbrace{\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\cdot\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})}_{\text{direction to }\mathbf{x}_{t}\text{}}+\underbrace{\sigma_{t}\boldsymbol{\epsilon}_{t}}_{\text{noise}},$ (3)
where $\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. Setting $\sigma_{t}=0$ yields a deterministic sampling process known as DDIM [52], which defines an invertible trajectory between data and latent space. Following [24, 44, 12, 39], we adopt this DDIM formulation for counterfactual generation, as detailed below.
Abduction. We implement the abduction function $\mathbf{u}:=f^{-1}_{\theta}(\mathbf{x},\mathbf{pa})$ using the DDIM forward trajectory. Given an observed image $\mathbf{x}_{0}$ (i.e. $\mathbf{x}$ in $f^{-1}_{\theta}$) and conditioning vector $\mathbf{c}$ (the embedding of semantic parents $\mathbf{pa}$), the latent $\mathbf{x}_{T}$ serves as a deterministic estimate of the exogenous noise $\mathbf{u}$:
$\mathbf{x}_{t+1}=\sqrt{{\alpha}_{t+1}}\cdot\hat{\mathbf{x}}_{0}+\sqrt{1-{\alpha}_{t+1}}\cdot\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,\mathbf{c}),\quad t=0,\dots,T-1,$ (4)
where the clean estimate $\hat{\mathbf{x}}_{0}$ at each step is computed as:
$\hat{\mathbf{x}}_{0}=\frac{1}{\sqrt{{\alpha}_{t}}}\left(\mathbf{x}_{t}-\sqrt{1-{\alpha}_{t}}\cdot\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})\right),$ (5)
Action. We apply an intervention to the semantic attributes $\mathbf{pa}$ (e.g., do(Male = 1)), and propagate the effect through the causal graph using invertible flows as in [36, 9]. This yields the counterfactual attribute vector $\widetilde{\mathbf{pa}}$ and its embedding $\tilde{\mathbf{c}}$.
Prediction. We implement the structural assignment $\tilde{\mathbf{x}}:=f_{\theta}(\mathbf{u},\widetilde{\mathbf{pa}})$ under the modified condition $\tilde{\mathbf{c}}$ using the DDIM reverse trajectory, with $\mathbf{u}=\mathbf{x}_{T}$ the exogenous noise estimated in eq. 4:
$\mathbf{x}_{t-1}=\sqrt{{\alpha}_{t-1}}\cdot\hat{\mathbf{x}}_{0}+\sqrt{1-{\alpha}_{t-1}}\cdot\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,\tilde{\mathbf{c}}),\quad t=T,\dots,1,$ (6)
where $\hat{\mathbf{x}}_{0}$ is computed as in eq. 5. The final output $\tilde{\mathbf{x}}_{0}$ is the predicted counterfactual $\tilde{\mathbf{x}}$.
In practice, decoding under the counterfactual condition $\tilde{\mathbf{c}}$ using the conditional denoiser alone may be insufficient for producing effective counterfactuals. Additional guidance is often required to steer generation toward the desired intervention [44, 45, 24, 12, 58, 53, 39, 25]. In this work, we adopt classifier-free guidance (CFG) to enhance counterfactual fidelity and alignment with the specified intervention.
2.3 Classifier-Free Guidance
CFG [16] is a widely adopted technique in conditional diffusion models. It enables conditional generation without requiring an external classifier by training a single denoising model to operate in both conditional and unconditional modes. During training, the model learns both $p_{\theta}(\mathbf{x}_{t}\mid\mathbf{c})$ and $p_{\theta}(\mathbf{x}_{t}\mid\varnothing)$ by randomly replacing $\mathbf{c}$ with a null token $\varnothing$.
At inference time, CFG samples from a sharpened posterior:
$p^{\omega}(\mathbf{x}_{t}\mid\mathbf{c})\propto p(\mathbf{x}_{t})\cdot p(\mathbf{c}\mid\mathbf{x}_{t})^{\omega},$ (7)
where $\omega\geq 0$ is a guidance weight controlling the strength of conditioning. This corresponds to interpolating between the unconditional and conditional scores:
$\nabla_{\mathbf{x}_{t}}\log p^{\omega}(\mathbf{x}_{t}\mid\mathbf{c})=\nabla\log p(\mathbf{x}_{t})+\omega\cdot\left(\nabla\log p(\mathbf{x}_{t}\mid\mathbf{c})-\nabla\log p(\mathbf{x}_{t})\right).$ (8)
In practice, this is implemented by interpolating denoised predictions:
@@ 2.3 Classifier-Free Guidance
$\boldsymbol{\epsilon}_{\text{CFG}}(\mathbf{x}_{t},t,\mathbf{c})=\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t\mid\varnothing)+\omega\cdot\left(\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})-\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,\varnothing)\right).$ (9)
With CFG, abduction is the same as in eq. 4, and action remains unchanged. The only difference lies in prediction, where the conditional denoiser is replaced with the guided score $\boldsymbol{\epsilon}_{\mathrm{CFG}}(\mathbf{x}_{t},t,\tilde{\mathbf{c}})$ to enhance counterfactual effectiveness [45, 24]:
@@
$\hat{\mathbf{x}}_{0}$ $=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\cdot\boldsymbol{\epsilon}_{\text{CFG}}(\mathbf{x}_{t},t,\tilde{\mathbf{c}})\right),$ (10)
@@
$\mathbf{x}_{t-1}$ $=\sqrt{\alpha_{t-1}}\cdot\hat{\mathbf{x}}_{0}+\sqrt{1-\alpha_{t-1}}\cdot\boldsymbol{\epsilon}_{\text{CFG}}(\mathbf{x}_{t},t,\tilde{\mathbf{c}}).$ (11)
Despite its effectiveness, standard CFG applies a single global guidance weight $\omega$ uniformly across the entire counterfactual embedding $\tilde{\mathbf{c}}$, which typically encodes multiple attributes—some of which may not have been altered during the intervention. In counterfactual generation, however, only a subset of attributes in $\tilde{\mathbf{c}}$ (i.e., those affected by the intervention) should be emphasized, while the remaining attributes should remain invariant. Applying the same guidance strength to all elements of $\tilde{\mathbf{c}}$ violates this principle, and can cause unintended changes to invariant attributes. This misalignment is called attribute amplification [63], which violates the causal relationship as pre-defined in structural causal models and undermines counterfactual soundness [32].
To address these limitations, we propose a structured alternative that assigns separate guidance weights to semantically or causally defined groups of attributes.
3 Method
In this section, we present our Decoupled Classifier-Free Guidance (DCFG) for counterfactual image generation. We begin by introducing an attribute-split conditioning embedder (section 3.1), which explicitly separates individual attributes in the embedding space to enable selective control. Building on this, we then describe our DCFG formulation (section 3.2), which allows distinct guidance strengths to be applied to different subsets of attributes. Finally, we present how DCFG is integrated into DDIM-based counterfactual inference (section 3.3), detailing its application across abduction, action, and prediction steps using causally defined attribute groupings.
3.1 Attribute-Split Conditioning Embedding
In practice, raw conditioning inputs such as discrete labels or structured attributes (e.g., a patient’s sex, race, disease status) are not used directly in diffusion models but transformed into dense vectors using embedding functions—typically via small MLPs [10], convolutional encoders [68], or transformer-based text encoders [16, 42]. These embeddings align semantic or categorical inputs with the model’s internal representation space, but conventional designs often entangle multiple attributes into a single vector, making it difficult to independently control them during sampling.
To address this, we introduce an attribute-split conditioning embedding that preserves the identity of each attribute in the embedding space. Let $pa_{i}$ denote the raw value of the $i$-th parent attribute (e.g., a binary indicator or scalar). Each $pa_{i}$ is embedded independently via a dedicated MLP $\mathcal{E}_{i}:\mathbb{R}^{d_{i}}\rightarrow\mathbb{R}^{d}$, and the final condition vector is formed by concatenating the outputs:
$\mathbf{c}=\operatorname{concat}\left(\mathcal{E}_{1}(pa_{1}),\dots,\mathcal{E}_{K}(pa_{K})\right),\quad\mathbf{c}\in\mathbb{R}^{K\cdot d}.$ (12)
This architecture provides a flexible representation where each attribute is explicitly disentangled at the embedding level. As a result, we can selectively null-tokenize or modulate individual attributes at inference time—enabling fine-grained control. Throughout the rest of the paper, we denote the semantic attribute vector as $\mathbf{pa}$ and the corresponding embedding vector as $\mathbf{c}$, as defined in eq. 12.
3.2 Decoupled Classifier-Free Guidance
To overcome the limitations of CFG and enable more precise, causally aligned control in counterfactual image generation, we propose Decoupled Classifier-Free Guidance (DCFG). Rather than applying a single guidance weight uniformly to the entire conditioning vector, we partition semantic attributes $\mathbf{pa}$ into $M$ disjoint groups $\mathbf{pa}^{(1)},\dots,\mathbf{pa}^{(M)}$, and apply a separate guidance weight $\omega_{m}$ to each group. Let $\mathbf{pa}=(pa_{1},\dots,pa_{K})$ denote the vector of semantic parent attributes. Under the assumption that the groups $\mathbf{pa}^{(1)},\dots,\mathbf{pa}^{(M)}$ are conditionally independent given the latent variable $\mathbf{x}_{t}$, we obtain the following factorized proxy posterior:
Proposition 1 (Proxy Posterior for DCFG).
Under the assumption $p(\mathbf{pa}\mid\mathbf{x}_{t})=\prod_{m=1}^{M}p(\mathbf{pa}^{(m)}\mid\mathbf{x}_{t})$, the sharpened proxy posterior becomes:
$p^{\omega}(\mathbf{x}_{t}\mid\mathbf{pa})\propto p(\mathbf{x}_{t})\cdot\prod_{m=1}^{M}p(\mathbf{pa}^{(m)}\mid\mathbf{x}_{t})^{\omega_{m}},$ (13)
where $\omega_{m}\geq 0$ controls the guidance strength for group $m$.
A full derivation and gradient-based justification for this proxy posterior are provided in Appendix B. The corresponding gradient used in score-based diffusion sampling is:
$\nabla_{\mathbf{x}_{t}}\log p^{\omega}(\mathbf{x}_{t}\mid\mathbf{pa})=\nabla\log p(\mathbf{x}_{t})+\sum_{m=1}^{M}\omega_{m}\cdot\left(\nabla\log p(\mathbf{x}_{t}\mid\mathbf{pa}^{(m)})-\nabla\log p(\mathbf{x}_{t})\right).$ (14)

### v5

# Factored Classifier-Free Guidance

@@ Abstract
Counterfactual generation aims to simulate realistic hypothetical outcomes under causal interventions. Diffusion models have emerged as a powerful tool for this task, combining DDIM inversion with conditional generation and classifier-free guidance (CFG). In this work, we identify a key limitation of CFG for counterfactual generation: it prescribes a global guidance scale for all attributes, leading to significant spurious changes in inferred counterfactuals. To mitigate this, we propose Factored Classifier-Free Guidance (FCFG), a flexible and model-agnostic guidance technique that enables attribute-wise control following a causal graph. FCFG complements recent advances in classifier-free guidance and can be seamlessly extended to advanced guidance schemes such as CFG++ and APG. Our experiments demonstrate that FCFG significantly improves the axiomatic soundness of inferred counterfactuals across both natural and medical image datasets, mitigating spurious amplification effects, and enhancing counterfactual reversibility.
Machine Learning, Diffusion Models, Counterfactual Generation, Classifier-Free Guidance, Causal Representation Learning
@@ 1 Introduction
Counterfactual generation is considered to be fundamental to causal reasoning (Peters et al., 2017; Bareinboim et al., 2022), allowing us to explore hypothetical scenarios such as: ‘How would this patient’s disease have progressed if they had received treatment A instead of treatment B?’. Answering such causal questions is important across various domains, such as healthcare (Castro et al., 2020), fairness (Kusner et al., 2017) and scientific discovery (Narayanaswamy et al., 2020). There has been a growing interest in generating counterfactual images using deep generative models, aiming to simulate how visual data would change under hypothetical interventions. Recent works build Structural Causal Models (SCMs) (Pearl, 2009) using deep generative model components such as normalizing flows (Rezende and Mohamed, 2015), Variational Autoencoders (VAEs) (Kingma and Welling, 2013; Child, 2020) and diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020), enabling principled counterfactual inferences via abduction-action-prediction (Pawlowski et al., 2020; Sanchez and Tsaftaris, 2022; Ribeiro et al., 2023; Wu et al., 2025; Rasal et al., 2025; Ribeiro et al., 2025a).
@@
Diffusion models have emerged as the state-of-the-art approach for image synthesis, achieving unprecedented fidelity and perceptual quality (Dhariwal and Nichol, 2021; Podell et al., 2023). Many previous works have explored diffusion models for counterfactual generation (Sanchez et al., 2022b, a; Pérez-García et al., 2024; Kumar et al., 2025; Rasal et al., 2025; Ribeiro et al., 2025b), leveraging Denoising Diffusion Implicit Models (DDIM) (Song et al., 2020) to deterministically encode images into a latent space, followed by conditional generation with modified attributes. Conditioning is typically enforced through guidance mechanisms, either with external classifier scores (Dhariwal and Nichol, 2021) or through Classifier-Free Guidance (CFG) (Ho and Salimans, 2022), which combines conditional and unconditional score estimates using a scalar guidance weight $\omega$. Combining DDIM inversion and guided conditional decoding has also become the dominant paradigm in diffusion-based image editing (Couairon et al., 2022; Wallace et al., 2023; Hertz et al., 2022; Epstein et al., 2023).
In counterfactual generation, CFG is often used to improve intervention effectiveness, by making the generated sample more strongly reflect the intended attribute change. While recent works have proposed refinements to classifier-free guidance to improve sampling fidelity (Chung et al., 2025; Sadat et al., 2025), both standard CFG and these extensions rely on a single global guidance signal applied to the entire conditioning embedding. In this work, we identify that in diffusion-based counterfactual generation, this uniform strengthening can inadvertently amplify attributes that should remain stable under a given causal intervention. We refer to this failure mode as attribute amplification, following the terminology introduced by Xia et al. (2024).
While Xia et al. (2024) observed attribute amplification in previous counterfactual models due to predictor-based fine-tuning (Ribeiro et al., 2023), we identify a distinct CFG-induced failure mode: even when a causal graph or semantic grouping is known, a single global guidance weight can amplify intervention-relevant and intervention-invariant attributes together. This behavior is induced by the guidance mechanism itself, rather than merely by dataset artefacts or causal-graph misspecification. It can violate the intended causal graph by modifying attributes outside the causal pathway and can also cause the generation trajectory to drift from the original data manifold, degrading identity preservation (Mokady et al., 2023). Thus, addressing CFG-induced attribute amplification is critical for the reliable use of guided diffusion models in counterfactual inference.
@@
To address the spurious effects of CFG under causal interventions, we propose Factored Classifier-Free Guidance (FCFG), a general inference-time guidance technique that significantly mitigates attribute amplification without requiring any changes to the underlying diffusion model. Unlike standard CFG, FCFG assigns separate weights to attribute groups, allowing for fine-grained, interpretable control over the generative process. In contrast to compositional diffusion approaches that rely on spatial masking (Shen et al., 2024) or multiple conditional models (Liu et al., 2022), FCFG operates at the semantic attribute level using a single diffusion model. The FCFG framework is general and supports arbitrary partitions of semantic attributes under mild independence assumptions. For counterfactual generation, we instantiate FCFG by grouping attributes according to their causal roles (e.g., intervened vs. invariant) and applying distinct guidance to each group. This assumes access to structured conditioning variables and a pre-specified causal or semantic grouping; FCFG does not address the upstream problem of causal discovery. Crucially, by factoring guidance and focusing it on the intended intervention, FCFG reduces the risk of the generation trajectory drifting away from the original data manifold (Yang et al., 2023; Mokady et al., 2023; Tang et al., 2024). Additionally, FCFG modifies only how the conditioning signal is factored and weighted, enabling its integration with standard CFG as well as recent guidance refinements such as CFG++ (Chung et al., 2025) and APG (Sadat et al., 2025). The contributions of this work are the following:
  1. 1.
    We identify and analyze the problem of attribute amplification in CFG, where a single global guidance scale indiscriminately strengthens all conditioning attributes, causing spurious changes in attributes that should remain invariant under a given intervention;
  2. 2.
@@
    We propose Factored Classifier-Free Guidance (FCFG), a simple, flexible, and model-agnostic extension of CFG that assigns separate guidance weights to attribute groups and supports arbitrary groupings at inference time under mild independence assumptions;
  3. 3.
    Through extensive experiments on challenging real-world datasets, including medical imaging, we show that FCFG reduces spurious effects, improves intervention effectiveness, and enhances counterfactual reversibility, and is compatible with recent guidance refinements such as CFG++ and APG.
@@ 2 Background
Structural Causal Models.
SCMs (Pearl, 2009) consist of a triplet $\langle U,A,F\rangle$, where $U{=}\{u_{i}\}_{i=1}^{K}$ denotes the set of exogenous variables, $A{=}\{a_{i}\}_{i=1}^{K}$ the set of endogenous variables, and $F{=}\{f_{i}\}_{i=1}^{K}$ a collection of structural assignments such that each variable $a_{k}$ is determined by a function $f_{k}$ of its parents $\mathbf{pa}_{k}{\subseteq}A{\setminus}a_{k}$ and its corresponding noise $u_{k}$, such that $a_{k}{\coloneqq}f_{k}(\mathbf{pa}_{k},u_{k})$. SCMs enable causal reasoning and interventions via the do-operator, e.g., setting a variable $a_{k}$ to a fixed value $c$ through $do(a_{k}{\coloneqq}c)$. In this work, we focus on generating image counterfactuals and implement the underlying image synthesis mechanism using diffusion models. We assume access to structured semantic parent variables and a pre-specified causal graph or semantic grouping over them; learning this graph is an orthogonal causal discovery problem and is outside the scope of this work.
Counterfactual inference.
@@
2.1 Diffusion Models for Counterfactual Inference
Diffusion models (DMs) (Sohl-Dickstein et al., 2015; Ho et al., 2020) are latent variable models designed to generate data by gradually removing Gaussian noise from $\mathbf{x}_{T}{\sim}\mathcal{N}(\mathbf{0},\mathbf{I})$ over $T$ steps. Given a clean data sample $\mathbf{x}_{0}{\sim}p_{\text{data}}$, the forward noising process is defined as follows:
$\mathbf{x}_{t}=\sqrt{\alpha_{t}}\,\mathbf{x}_{0}+\sqrt{1-\alpha_{t}}\,\boldsymbol{\epsilon},\qquad\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (1)
where $\{\alpha_{t}\}_{t=0}^{T}$ is a chosen noise schedule with $\alpha_{t}{\in}(0,1]$, $\alpha_{0}{=}1$ and $\alpha_{T}{\approx}0$. To learn the reverse process, a parameterized network $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})$ is trained to predict the added noise from noisy inputs. We adopt the conditional diffusion model formulation, where $\mathbf{c}$ denotes an embedding of semantic parent attributes $\mathbf{pa}$ used as conditioning. In our implementation, $\mathbf{c}$ is formed from attribute-specific MLP embedders that are trained jointly with the denoising network end-to-end, rather than as independently pretrained feature extractors. The training objective minimizes the noise prediction loss:
@@ 2.1 Diffusion Models for Counterfactual Inference
$\underset{\theta}{\min}\;\mathbb{E}_{\mathbf{x}_{0}\sim p_{\text{data}},\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),t\sim p_{t}}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})\right\|^{2}_{2}\right].$ (2)
At inference time, data samples are generated by progressively denoising $\mathbf{x}_{T}$ from time $T$ to time 0. Following Song et al. (2020), the denoising step from $\mathbf{x}_{t}$ to $\mathbf{x}_{t-1}$ is given by:
@@
$\mathbf{x}_{t-1}$ $=\sqrt{\alpha_{t-1}}\left(\frac{\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\,\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})}{\sqrt{\alpha_{t}}}\right)$ (3)
@@
$\quad+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\,\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})+\sigma_{t}\,\boldsymbol{\epsilon}_{t},$
where $\boldsymbol{\epsilon}_{t}{\sim}\mathcal{N}(\mathbf{0},\mathbf{I})$. Setting $\sigma_{t}{=}0$ yields a deterministic sampling process known as DDIM (Song et al., 2020), which defines an invertible trajectory between data and latent space. Following Sanchez and Tsaftaris (2022); Sanchez et al. (2022a); Fontanella et al. (2024); Pérez-García et al. (2024); Rasal et al. (2025), we adopt this DDIM formulation for counterfactual generation, as summarized below; full algorithmic details are provided in App. C.
Abduction. We implement the abduction function $\mathbf{u}{=}f^{-1}_{\theta}(\mathbf{x}_{0},\mathbf{pa})$ using the DDIM forward trajectory. Given an observed image $\mathbf{x}_{0}$ and conditioning $\mathbf{c}$ representing an embedding vector of semantic parents $\mathbf{pa}$, the latent $\mathbf{x}_{T}$ serves as a deterministic estimate of the exogenous noise $\mathbf{u}$:
$\mathbf{x}_{t+1}$ $=\sqrt{\alpha_{t+1}}\,\hat{\mathbf{x}}_{0}+\sqrt{1-\alpha_{t+1}}\,\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,\mathbf{c}),$ (4)
$\hat{\mathbf{x}}_{0}$ $=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\,\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})\right),$
for $t=0,{\dots},T-1$, where $\hat{\mathbf{x}}_{0}$ denotes the model’s estimate of the clean image at time $t$. The full abduction procedure is summarized in Algorithm 1.
Action. We apply an intervention to the semantic attributes $\mathbf{pa}$, e.g., do(Male=1), and propagate the effect through the causal graph using invertible flows as in Pawlowski et al. (2020); Ribeiro et al. (2023). This yields the counterfactual attribute vector $\widetilde{\mathbf{pa}}$ and its embedding $\tilde{\mathbf{c}}$. The full action procedure is summarized in Algorithm 2.